### v1

# Predicting Maintenance Cessation of Open Source Software Repositories with An Integrated Feature Framework

Abstract.
The maintenance risks of open source software (OSS) projects pose significant threats to the quality, security, and resilience of modern software supply chains. While prior research has proposed diverse approaches for predicting OSS maintenance risk—leveraging signals ranging from surface features (e.g., stars, commits) to social network analyses and behavioral patterns—existing methods often suffer from ambiguous operational definitions, limited interpretability, and datasets of insufficient scale or generalizability. In this work, we introduce “maintenance cessation”, grounded in both explicit archival status and rigorous semantic analysis of project documentation. Building on this foundation, we curate a large-scale, longitudinal dataset of 115,466 GitHub repositories—encompassing 57,733 confirmed cessation events—complemented by comprehensive, timeline-based behavioral features. We propose an integrated, multi-perspective feature framework for predicting maintenance cessation, systematically combining user-centric features, maintainer-centric features and project evolution features. AFT survival analysis demonstrates a high C-index (0.846), substantially outperforming models relying only on surface features. Feature ablation and SHAP analysis further confirm the effectiveness and interpretability of our approach. Finally, we demonstrate real-world applicability by deploying a GBSA classifier in the openEuler ecosystem for proactive package risk screening. Our work establishes a scalable, interpretable foundation for maintenance-risk prediction, enabling reproducible risk management across large-scale open source ecosystems.
1. Introduction
Open-source software (OSS) is the foundation of modern software engineering, powering diverse applications and enterprise systems. Collaborative platforms such as GitHub have enabled developers to reuse third-party OSS components at an unprecedented scale, significantly accelerating development and reducing cost. According to the 2025 Open Source Security and Risk Analysis (OSSRA) report (Synopsys, 2025), over 97% of analyzed software projects across industries contain OSS dependencies, underscoring the centrality of OSS in contemporary software supply chains.
Meanwhile, the growing dependence on open source software means the growing risk of ecosystem-wide disruptions. The halt of maintenance is not rare in the open source ecosystem, and it poses the sudden risk of software defects, security vulnerabilities, and more to numerous downstream users and dependent software systems (Kula et al., 2017). The OSSRA report finds that 91% of audited OSS components exhibited no clear signs of maintenance in the past two years (Synopsys, 2025), highlighting the prevalence of this issue. And notable, high-profile OSS projects may also run into discontinuation—well-known examples include Atom (Atom, [n. d.]), Brackets (adobe, [n. d.]), and faker.js (Marak, [n. d.]).
Proactive prediction of unmaintainance in OSS projects is a promising approach to mitigate this risk. By warning developers and organizations about the unmaintenance risk, it enables them to take proactive countermeasures including vendoring, removing, or migrating risky dependencies to anticipate disruptions. While extensive research has explored the potential of data-driven unmaintenance risk prediction (Samoladas et al., 2010a; Xia et al., 2022), two gaps must be addressed for the prediction techniques to be practical.
The first gap lies in the definition and identification of unmaintenance. Many of previous studies of project “deprecation” or “failure” defined unmaintenance based on the activity features, they often assume a project is abandoned after a fixed threshold of inactivity (Coelho and Valente, 2017; Liao et al., 2023; Robinson et al., 2022; Malviya Thakur et al., 2025). However, we found counter-examples of “revival” in real-life open source projects to their assumption (Section 3.1), indicating that inactivity is not a reliable proxy of project unmaintainance.
The second gap is in the prediction techniques. Existing approaches for prediction typically use surface features such as the number of star, commit, or pull request (PR), which are volatile, susceptible to manipulation (He et al., 2024), and insufficient for capturing deeper sustainability dynamics such as governance transitions, contributor attrition, or changes in user engagement. Other approaches that incorporate more high-level features and advanced methods often lack semantic interpretability, which questions their practical reliability and generalizability.
In this paper, we introduce the term maintenance cessation as a clear and definitive definition of the end of project maintenance. Specifically, we identify maintenance cessation based on the union of two criteria: (1) explicit GitHub archived status (GitHub Documentation, 2025; GitHub, nd), and (2) unambiguous maintenance cessation statements in project documentation (Coelho and Valente, 2017). Cessation is both semantically precise and task-specific, and maintenance cessation more closely captures the genuine discontinuation of maintenance, explicitly excludes projects that are merely “quiet” or low-activity but remain responsibly maintained (Naberezny et al., 2011; Ashkenas et al., 2013; Suuronen et al., 2022).
Based on the definition of maintenance cessation, we curate a large-scale longitudinal dataset of 115,466 GitHub repositories, including 57,733 with validated maintenance cessation events and comprehensive behavioral timelines (Section 3.3). We then propose a multi-dimensional, interpretable feature framework that goes beyond surface features by jointly modeling: (1) user influence using a time-aware, weighted bipartite user-project interaction network; (2) maintainer behavior, including activity recency and response latency; and (3) project evolution characteristics, such as community role balance and the shift between feature and bugfix contributions (Section 4.1). Using an accelerated failure time (AFT) survival model, we empirically demonstrate that our approach substantially outperforms existing methods, while providing improved interpretability and early warning. We further train a GBSA classifier and integrated it into openEuler’s risk assessment pipeline (Section 5). It proves that our framework is practical in assisting the risk control and management of real-life large-scale software supply chains.
To sum up, in this paper, we:
  • Proposed a precise definition of OSS maintenance cessation.
  • Constructed a large-scale, longitudinally labeled dataset of OSS repositories.
  • Designed a multi-dimensional, interpretable feature framework for predicting maintenance cessation.
  • Empirically demonstrated substantial predictive improvements over existing methods.
  • Validated the real-world utility of our approach through deployment in the openEuler ecosystem.
We believe our approach lays a crucial foundation for the future of software supply chain risk assessment, OSS dependency management, and sustainable ecosystem governance.
2. Background
The challenge of predicting when an open source project will cease to be maintained—whether referenced as “abandonment,” “deprecation”,“failure”, or more broadly as sustainability—has generated extensive research attention in the software engineering domain. Early work in this area primarily adopted statistical survival analysis approaches drawn from reliability engineering to model project longevity and risk. For example, models such as the Cox proportional hazards model and the Kaplan-Meier estimator were used to investigate how factors like the number of active contributors and the presence of major version releases correlated with continued OSS project maintenance or survival (Samoladas et al., 2010a; Robinson et al., 2022; Khondhu et al., 2013; Samoladas et al., 2010b; Li et al., 2022a; Valiev et al., 2018; Mujahid et al., 2022).
With the increasing availability of large-scale data from social coding platforms (e.g., GitHub), more recent research has turned to machine learning (ML)-based methods. Most existing approaches operationalize “abandonment” or “deprecation” based on symptoms such as prolonged inactivity in commits, issues, or PR, and less often, declarations from maintainers in documentation (Liao et al., 2023; Moid et al., 2021; Li et al., 2022b). Feature engineering in these works has focused primarily on surface-level, readily available project activity statistics—counts of stars, commits, issues, PRs, and sometimes doc structure or languages. Some research has explored augmenting these signals with social or network-derived features, such as developer influence or collaboration centrality measures, or has employed more advanced architectures including ensemble boosting and graph neural networks (Ren et al., 2020; Liao et al., 2023; Xia et al., 2022; Coelho et al., 2018, 2020).
Despite these advances, several key limitations hamper the robustness, interpretability, and practical adoption of the current state-of-the-art. Specifically, we identify several major areas where existing research falls short:
Ambiguous or oversimplified definitions of maintenance cessation. The majority of prior work equates “abandonment” or “deprecation” with simple inactivity over a fixed period—typically time since last commit or last issue. However, these terms are ambiguous in meaning, and ignores the reality that many mature projects may operate with infrequent updates but still provide essential maintenance and user support, while some genuinely unmaintained projects might not be captured by activity alone. The lack of a universally precise and robust definition for true maintenance cessation.
Over-reliance on surface features. Most prior approaches rely primarily on surface features (e.g., stars, commits, PRs, issues), which, despite being easy to obtain, may fail to capture deeper signals of project sustainability. The limitations of such surface features are further discussed in Section 4.1.1.
Opaque feature semantics in advanced models. Some studies attempt to improve prediction by extending surface features with network-based or behavioral signals, or by employing complex models (Ren et al., 2020; Liao et al., 2023; Xia et al., 2022). However, these approaches often yield limited predictive performance and lack interpretable explanations of how such features contribute to outcomes, relying instead on model capacity alone. This opacity limits their practical value to ecosystem maintainers who need transparent, trustworthy insights for decision-making.
Limited data scale and ecosystem coverage. Existing datasets are often small, highly curated, or biased towards repositories with high visibility (e.g., high star counts), and as such do not represent the diversity of OSS projects. As a consequence, findings from these studies may lack generalizability to the broader software ecosystem.
Insufficient empirical validation and practical deployment. Even as machine learning techniques become more sophisticated, their practical effectiveness is rarely examined at ecosystem scale. Most approaches are evaluated only in isolated, research-focused settings and are not validated within actual package management or release governance workflows.
Recent advances have begun to address some of these shortcomings, but major gaps persist, particularly in integrating behavioral signals, scaling to large real-world ecosystems, and providing definitions and methods that are both theoretically sound and operationally actionable.
3. Methodology
This section presents our unified methodological framework for predicting maintenance cessation of OSS repositories. Our approach encompasses formal problem definition, a scalable and rigorous pipeline for data construction and labeling, the design of interpretable and discriminative feature sets characterizing user, maintainer, and project dynamics, and robust predictive modeling strategies tailored to both research and operational deployment.
3.1. Problem Definition
Accurately predicting when an OSS project will cease to be actively maintained is a critical but challenging task for both researchers and practitioners. However, before prediction is possible, a precise and actionable definition of “maintenance cessation” is essential. Existing literature uses various terms such as “abandonment”, “deprecation” or “failure”, but often with vague meaning and inconsistent operationalization: some studies rely solely on indicators of prolonged inactivity (e.g., the absence of commits within a fixed period) (Coelho and Valente, 2017; Liao et al., 2023; Robinson et al., 2022; Malviya Thakur et al., 2025), while others require declarations from maintainers in documentation or metadata. Moreover, these terms carry domain-specific meanings in other areas of software engineering, which can lead to semantic confusion (Robbes et al., 2012; Haryono et al., 2021; Song et al., 2024). This ambiguity causes significant challenges for scientific reproducibility and practical application:
Over-reliance on (in)activity: Many mature OSS projects exhibit infrequent updates but remain under responsible maintenance. For example, supervisor (Naberezny et al., 2011), backbone (Ashkenas et al., 2013), and underscore.string (Suuronen et al., 2022) may go several months—or even over a year—without any commits. However, their maintainers still assume stewardship over the repositories, occasionally performing code modifications when critical issues arise. Treating such low-frequency activity as a sign of unmaintained can result in high false-positive rates.
Semantic confusion: Terms such as “abandonment,” “deprecation,” or “failure” may refer to ceasing feature development, stopping issue triage, discontinuing support for older versions, or ending all development and support. Without explicit definition and consistent definition, label quality and downstream prediction suffer.
In this work, we adopt the term maintenance cessation to describe the precise condition in which the maintainers have explicitly and definitively declared the end of maintenance for a software repository. This term is deliberately chosen not merely to avoid the ambiguity of commonly used alternatives like abandonment or deprecation, but because it clearly and unambiguously communicates the specific event we aim to detect-namely, an intentional, declared termination of maintenance. Compared to other vague or overloaded terms, maintenance cessation is immediately interpretable and directly aligned with our task objective. It emphasizes the cessation of active stewardship, without presuming disuse, obsolescence, or removal from the ecosystem. Our definition explicitly excludes projects that are simply low-activity, stable, or “quiet” but are still under responsible maintenance.
Accordingly, we identify maintenance cessation as occurring if, and only if, at least one of the following two conditions is satisfied:
  1. (1)
    Explicit archival status: The repository is marked as “archived” on GitHub, which irreversibly freezes the repository in a read-only mode and reflects a maintainer’s explicit cessation of maintenance (GitHub, nd; GitHub Documentation, 2025). The timestamp of archiving is recorded as the cessation event.
  2. (2)
    Semantic cessation statement: The repository’s top-level documentation (e.g., README, project description) contains direct, unambiguous textual statements indicating that the repository is deprecated, no longer maintained, unmaintained, or has reached end-of-life. To minimize ambiguity, we detect such statements through a hybrid pipeline that combines targeted keyword filtering, manual verification of a large sample, and large language model (LLM) and transformer-based classification, as detailed in Section 3.3.
This dual-criteria definition ensures: (1) precision—capturing only verifiable maintenance cessation events, not mere inactivity, and (2) recall—identifying cases explicitly signaled in text but not only using GitHub’s archival feature. These clear labels form the basis for predictive modeling in the remainder of this work.
3.2. Methodology Overview
Figure 1 provides an overview of our methodological framework for predicting maintenance cessation of OSS repositories. Our approach is structured around five core components, each tightly integrated to ensure scientific rigor, feature interpretability, and real-world applicability.
Figure 1. Methodology Overview.
  1. (1)
    Defining Maintenance Cessation. We establish a dual-criteria definition where repositories cease maintenance if explicitly archived or containing semantic cessation declarations.
  2. (2)
    Data Construction. We collect GitHub repositories (stars $\geq$32) and assign cessation labels through hybrid annotation combining keyword filtering, manual validation, and automated classification.
  3. (3)
    Feature Engineering. We develop a multi-perspective framework including surface features, user influence scores from bipartite networks, maintainer behavioral signals, and project evolution indicators.
  4. (4)
    Feature Effectiveness Validation. We validate features through survival analysis using AFT models with ablation studies and SHAP interpretability, supplemented by temporal case analyses.
  5. (5)
    Real-World Deployment. We implement risk prediction in openEuler ecosystem via GBSA classifiers applied to upstream package mappings.
By integrating precise definitions, scaleable data processing, interpretable feature frameworks, rigorous modeling, and operational validation—as visually summarized in Figure 1—our methodology delivers a reliable, reproducible, and actionable blueprint for large-scale OSS maintenance cessation prediction.
3.3. Data Construction and Preprocessing
To support robust, generalizable prediction and facilitate downstream research, we constructed a comprehensive, high-quality dataset capturing both longitudinal behavioral traces and rigorously validated cessation labels for a large population of OSS repositories. This process comprised repository selection, multi-source data collection, careful labeling, temporal aggregation, and quality control. The Data Construction and Preprocessing box in Figure 1 shows the general process.
Repository Selection and Scope.
We target repositories hosted on GitHub (GitHub, 2007), the world’s largest collaborative development platform. To balance representativeness and tractability, we include all non-fork repositories with at least 32 stars, following both prior studies and API limit considerations. This threshold effectively filters out trivial or inactive repositories, while preserving coverage across diverse languages, domains, and activity levels. Our time span covers repositories created from 2011 through 2025, ensuring sufficient history for meaningful temporal analysis. Forked projects are excluded to avoid redundancy and ambiguous maintenance signals linked to upstream dependencies.
Ground Truth Labeling: Maintenance Cessation.
Accurately labeling the cessation of software maintenance poses a significant challenge due to its informal and distributed nature. We adopt a three-step pipeline to construct high-quality, reproducible ground truth labels for repository maintenance cessation:
  1. (1)
    Archived Repository Collection. We leverage the GitHub GraphQL API (Github, [n. d.]) to retrieve repositories explicitly marked as ”archived”. This status, once enabled by maintainers, freezes the repository in a read-only mode and is treated as a strong, platform-verified signal of maintenance cessation. To reduce noise from trivial or inactive projects, only repositories with more than 32 stars are retained, aligning with both practical and API efficiency considerations (Github, [n. d.]). The archival timestamp is recorded as the official cessation date.
  2. (2)
    Cessation Declaration Mining. As prior work (Coelho and Valente, 2017) has shown, many developers choose to declare “deprecation” or “abandonment” in README files or repository descriptions rather than use the archive flag. We identify such cases using a semantic statement extraction strategy: (1) A list of maintenance cessation keywords (e.g., “no longer maintained”, “deprecated”, “unmaintained”, etc.) is first used to retrieve candidate repositories. To mitigate false positives due to ambiguous usage (e.g., deprecating a dependency rather than the project itself), we perform manual annotation on 1,200 stratified samples. (2) The annotation task is conducted using Label Studio (HumanSignal, 2024), where two independent annotators label each sample based on both textual signals and repository activity metadata (e.g., commits, issues, PRs). The resulting labels achieve a Cohen’s kappa of 0.814, indicating substantial inter-annotator agreement. (3) Using 600 samples for training and 600 for validation, we fine-tune a SetFit classifier (Tunstall et al., 2022) based on sentence transformers (Reimers and Gurevych, 2019). Our best model, built on the paraphrase-mpnet-base-v2 checkpoint (sentence-transformers, 2022), achieves an accuracy of 0.96, recall of 0.96, and precision of 0.90.
  3. (3)
    LLM-Assisted Labeling. To scale high-precision labeling, we evaluated multiple large language models (LLMs) on the 1,200-sample benchmark. Among the tested models (including GPT-4o, DeepSeek variants, Claude 3.7, and Qwen series), GPT-4o demonstrated superior performance with accuracy, precision, recall, and F1 achieving 0.9463, 0.9429, 0.8985, and 0.9202. This model was consequently selected to label the remaining 37,065 samples. We observed 95.91% agreement between GPT-4o and our SetFit classifier. All repositories where GPT-4o and SetFit predictions disagreed were manually reviewed and annotated by authors to resolve discrepancies and ensure final label quality.
This process results in 57,733 repositories with confidently labeled maintenance cessation dates, derived from either archival flags or timestamped textual cessation declarations.
Longitudinal Behavioral Data Extraction.
To capture project evolution and enable feature engineering, we aggregate fine-grained, time-series records of repository activities using a combination of GHArchive event data (GHArchive, 2024) and the GitHub API. For each repository, we record (per month):
  • Code contributions: commits (including lines changed), PR creations.
  • Community engagement: issues, comments, stars.
  • Metadata changes: tags, description or README updates.
Compared with other public datasets (e.g., GHTorrent (Gousios, 2013)), GHArchive is leveraged for its completeness and efficiency in reconstructing historical events at scale. Events are parsed using distributed ETL pipelines and loaded into a column-oriented OLAP system (ClickHouse), supporting performant querying and flexible aggregation.
Figure 2. Maintenance lifespan distribution of ceased repositories.
Data Cleaning and Quality Control.
Quality control steps include: (1) Removal of repositories with inconsistent, corrupted, or incomplete event histories. (2) Cross-validation of label assignments to mitigate false positives, e.g., distinguishing actual project maintenance cessation from critical version maintenance cessation or temporary inactivity. (3) Statistical analysis of activity, star, and lifetime distributions for ongoing vs. ceased repositories, confirming representativeness and detecting potential annotation or sampling bias (see Table 1 and Figure 2).
Table 1. Distribution of star counts in the dataset.
Category Mean Std Min 50% Max
All Projects 382.88 2241.63 32 83 222257
Ceased Projects 268.79 443.33 32 126 2383
4. Feature Framework
Numerous factors contribute to the deaths of open source projects (Robbes et al., 2012; Sawant et al., 2016). We design a feature framework capturing multiple perspectives in OSS maintenance, including surface activity, user participation, maintainer behavior, and project evolution patterns. Our features are constructed from large-scale, longitudinal repository data, organized into three main categories: user-centric, maintainer-centric, and project evolution features. In the following subsections, we describe each group and its computation.
The remainder of this section systematically describes the design and computation of each feature category, and empirically justifies their inclusion via ablation and interpretability analysis in Section 4.2.
4.1. Feature Design
We first summarize the surface features commonly adopted in existing literature and then motivate our enhancements via new structured feature groups. Table 2 provides an overview of all major feature categories used in our framework.
Table 2. Categorized list of features used in the cessation prediction framework.
Feature Description
Surface Activity
stars Number of stars received by the repository.
commits Number of commits in the repository.
issues Number of issues opened in the repository.
prs Number of pull requests submitted.
tags Number of tags in the repository.
comments Number of comments made on issues and PRs.
User-Centric
weight User-centric feature (see Section 4.1.2).
weight_rank_pct Percentile rank of weight.
weight_zscore Z-score normalized value of weight.
Maintainer-Centric
latest_maintainer_activity_interval Time interval since the last maintainer action.
avg_response_time Average maintainer response time to issues and PRs.
response_decay_trend Trend of avg_response_time; positive indicates degradation.
Project Evolution (By default, data within the recent 6 months)
maintainer_contrib_ratio Proportion of maintainer activities over all repository activities.
contrib_diversity Contributor diversity measured by 1 - Gini index.
balance_index Structural balance combining maintainer and others contribution share.
activity_deviation Normalized deviation of recent 3-month activity from historical mean.
quarterly_deviation Deviation from periodic quarterly activity trends.
feature_ratio Ratio of feature-related PRs among all PRs.
bugfix_ratio Ratio of bug-fixing PRs among all PRs.
bugfix_feature_ratio Ratio of bugfix PRs to feature PRs, indicating maturity shift.
4.1.1. Surface Features
Prior work on OSS health and abandonment prediction has largely relied on surface-level features. These include cumulative or interval-based counts of stars, commits, issues, PRs, tags, and comments—all of which are directly accessible via public platform APIs and have intuitive appeal as signals of project activity or visibility (Liao et al., 2023; Moid et al., 2021; Robinson et al., 2022; Chen et al., 2014).
While these features provide a useful starting point, they exhibit several critical limitations in our setting:
Lack of user differentiation: Surfacer features treat all user actions equally, overlooking differences in user roles. For example, a star from a renowned developer is more meaningful than one from a novice, but surface features flatten such distinctions.
Vulnerability to manipulation: Surface features are easily inflated (e.g., via paid stars or bots (He et al., 2024)), undermining their reliability for predicting project health.
Limited expressive power: Surface features blur the line between maintainers and casual users, and cannot capture collaboration quality or project lifecycle changes—leading to missed signals about maintenance trajectories.
To address these shortcomings and move beyond the limitations of surface features, we next introduce three structured feature categories.
4.1.2. User-Centric Features
To overcome the limitations of surface activity features, which do not distinguish user roles and are prone to manipulation, we introduce user-centric features based on the structure and evolution of user-repository interactions. Our key insight is that users attracted to high-quality projects tend to be more experienced, and repositories that engage these users are more likely to be sustainable. Some studies have uncovered new information within networks (Clauset et al., 2005; Pinzger et al., 2008; Nagappan et al., 2010). We represent the OSS ecosystem as a weighted bipartite graph (Blincoe et al., 2016) (Figure 3) and iteratively propagate influence between users and repositories. This approach enables us to capture complex patterns of user influence and project health that surface features alone cannot reveal.
Figure 3. Illustration of the user-repository bipartite graph.
Interaction-Based Influence Modeling.
In our formulation, each user-repository pair is linked through various types of interactions, including starring, committing, forking, and issues. Building upon the intuition that the quality and timing of user engagement are more indicative of sustainability than sheer quantity, we simply assign interaction-specific base weights: $w_{orig}^{star}=1$, $w_{orig}^{commit}=8$, $w_{orig}^{fork}=4$, $w_{orig}^{issue}=2$. To further emphasize early or persistent engagement, a temporal decay factor is applied to each interaction, defined as:
(1) $p_{(up)}^{t}=\frac{k_{(up)}^{t}}{N_{p}},\quad D(p)=\frac{1}{1+p},\quad d_{(up)}^{t}=D(p_{(up)}^{t})$
[…]
Computation of Interaction Weights.
For each user-repository edge, interaction weights are computed as follows:
(2) $w_{(up)}^{star}=w_{orig}^{star}\times d_{(up)}^{t},\quad w_{(up)}^{fork}=\sum_{fork_{(up)}}w_{orig}^{fork}\times d_{(up)}^{t},$
$w_{(up)}^{issue}=\sum_{issue_{(up)}}w_{orig}^{issue}\times d_{(up)}^{t}$
To better quantify the technical contribution of commits, we additionally weight each commit by the logarithm of lines of code changed ($LOC$):
(3) $w_{commit}=\sum_{commits_{(up)}}w_{commit}^{orig}\times\lg LOC~\textbf{if}~LOC>0~\textbf{else}~0$
Total user-repository weights are then aggregated:
(4) $w_{(up)}=\sum_{o\in Set_{op}}w_{(up)}^{o},\quad{Set_{op}}=\{star,commit,fork,issue\}$
Propagation of Influence: HITS-Style Approach.
Inspired by HITS (Hyperlink-Induced Topic Search) (Kleinberg, [n. d.]; Prajapati, 2012), and leveraging the bipartite graph structure, we employ an iterative algorithm to propagate both influence and project quality as follows:
(5) $PQS(p)=\sum_{u\in U_{p}}w_{(up)}\times UIS(u),PQS(p)=\frac{PQS(p)}{\sum_{p^{{}^{\prime}}\in P}PQS(p^{{}^{\prime}})}$
(6) $UIS(u)=\sum_{p\in P_{u}}w_{(up)}\times PQS(p),UIS(u)=\frac{UIS(u)}{\sum_{u^{{}^{\prime}}\in U}UIS(u^{{}^{\prime}})}$
Here, $PQS(p)$ and $UIS(u)$ represent the project quality score for repository $p$ and the user influence score for user $u$, respectively. Scores are initialized to 1 and updated iteratively until convergence. Normalization ensures that scores are comparable among all repositories/users.
Implementation and Scalability.
Given the large scale of user-project activity on GitHub, scalable computation is essential. We implement this process using distributed Spark-based pipelines, operating on user-project interaction logs stored in TiDB (PingCAP Inc., 2017). This allows efficient monthly recalculation of influence features across 115,466 repositories and millions of users.
Robustness and Expressiveness.
This user-centric formulation offers several advantages. By taking into account the historical quality of user participation, it distinguishes habitual contributors from one-off actors and thus uncovers subtle signals of project health. Additionally, the graph-based design reduces susceptibility to manipulation, as artificially inflated surface features do not result in higher network-based influence. Finally, the incorporation of temporal weighting highlights sustained or early engagement, which is closely associated with long-term project sustainability.
Normalization for Temporal and Cross-Project Comparison.
Similar to many key features in software engineering (Goeminne and Mens, 2011; Zhang et al., 2021), our proposed user-centric features also exhibit a pronounced long-tail distribution. To ensure feature comparability across different time periods and repositories, we apply two normalization strategies: (1) percentile ranking, which transforms scores into relative ranks within each time window (Mujahid et al., 2022), and (2) Z-score normalization, with logarithmic transformation to address skewness:
(7) $w_{\%}=\frac{rank(w_{i})}{|i|},\quad w_{z}=\frac{\ln w_{i}-\mu}{\sigma}$
These normalized features are used in downstream models, ensuring stability and interpretability for survival analysis and classification.
4.1.3. Maintainer-Centric Features
Maintainers hold primary responsibility for project evolution, governance, and the formal declaration of maintenance cessation, their behavioral signals serve as salient predictors of project risk. To this end, we design maintainer-centric features that directly track maintainer engagement and responsiveness within the repository lifecycle.
Maintainer Identification.
We identify maintainers for each repository and time window by analyzing commit history and PR metadata. Specifically, a user is labeled as a maintainer if they either (1) perform direct commits, or (2) merge PRs. This operational criterion captures users with active governance roles, enabling accurate maintainer tracking over time.
Maintainer Activity Recency.
An important feature of governance continuity is the time elapsed since the last maintainer activity. For each repository and timestamp, we compute the maintainer inactivity interval as the duration since the most recent repository-level action performed by any identified maintainer. This inclusive approach captures the latest evidence of maintainer presence, beyond privileged operations like merges or direct commits. A prolonged inactivity interval strongly signals disengagement and is frequently a precursor to maintenance cessation. Compared to user activity features, this feature provides finer granularity in distinguishing actual maintenance cessation from superficial reductions in community activity.
Maintainer Response Latency.
Healthy open source projects typically exhibit prompt maintainer responses to externally submitted PRs and issues, indicating ongoing stewardship and community integration. As the risk of cessation grows, response times elongate or become absent altogether. We capture this dynamic through two related features:
  • Mean Response Time (Last 6 Months): The average delay for maintainers to address new PRs and issues within the most recent 6-month window, quantifying timeliness of governance actions.
  • Change in Response Time (Last 6 Months): The temporal trend, measured as the delta in mean response time over sequential intervals, providing early detection of deteriorating maintainer attention.
Contribution Ratio and Community Balance.
Stability in the proportion of maintainer involvement relative to overall project actions is a further signal of sustainable governance. Over-reliance on maintainers (or, conversely, abrupt decline in their participation) may indicate community imbalance or capacity exhaustion. We therefore include as features:
  • Maintainer Contribution Ratio: The fraction of all substantial actions (commits, merges, issues) performed by maintainers in a sliding window.
  • Imbalance Feature: Quantified as the squared deviation from an empirical ideal ratio (e.g., 0.5), emphasizing both under- and over-concentration of maintainer activity.
4.1.4. Project Evolution Features
Beyond user and maintainer dynamics, the long-term evolution of an OSS project reveals critical signals of sustainability, community health, and risk of maintenance cessation. To capture these deeper aspects, we construct a suite of project evolution features that model not only activity levels but also the changing collaboration structure, development focus, and periodic behavioral patterns over the lifecycle of a repository.
Community Participation Balance.
Mature open source projects typically follow an “onion model” comprising core maintainers, active contributors, and peripheral users. Disruption to the balance among these roles—such as disproportionate maintainer dominance or a decline in contributor diversity—often precedes project maintenance cessation. We operationalize this via:
  • Maintainer Contribution Ratio ($p$): The proportion of key repository interactions (e.g., commits, PR merges) performed by maintainers in a recent window (e.g., six months).
  • User Activity Gini Coefficient ($G$):
    (8) $G=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|}{2n^{2}\mu}$
    where $n$ is the number of users interacting with the project, $x_{i}$ the count of actions by user $i$, and $\mu$ the average number of actions. A high $G$ indicates concentration of activity and potential fragility in collaboration.
  • Participation Imbalance ($d$): The squared deviation from an ideal maintainer ratio: $d=(p-0.5)^{2}$. Lower values indicate a healthy, balanced distribution of responsibilities, while high values signal imbalance and increased cessation risk.
Development Focus via PR Categories.
The relative focus of project development—whether on adding new features or on bug fixing—naturally shifts as a repository matures. A transition from innovation-driven contributions to predominantly bug-fixing is often symptomatic of declining momentum: PRs are classified as feature-oriented or bug-fix based on curated keyword matching in PR titles (e.g., “feature”, “add” for features; “fix”, “bug” for bug-fixes). For each time window, we extract: the ratio of feature to bug-fix PRs, as well as the individual proportions of each category. A shrinking feature/bug-fix PR ratio is an early marker of functional stagnation and elevated risk of upcoming maintenance cessation.
Deviation from Periodic Activity.
Many OSS projects demonstrate regular, periodic activity patterns reflecting release planning or cyclical maintenance behaviors. Disruption or breakdown of these patternsoften signals the onset of decline:
  • Activity Deviation: For each repository, we compute the average of core activity features (e.g., commits, issues, PRs) over the recent quarter/3 months and compare this to their historical baseline using z-score normalization.
  • Quarterly Deviation: Activity in each recent quarter is standardized against the average of the previous three quarters, highlighting abrupt slowdowns or surges in development efforts.
A consistent pattern of negative deviation serves as a quantifiable warning signal preceding cessation events.
4.2. Feature Effectiveness Validation
Just as physiological indicators can predict health outcomes, behavioral traces of OSS repositories can serve as “biomarkers” for project vitality. In our setting, the interval from repository creation to maintenance cessation constitutes the software “lifespan.” Survival analysis is well suited for this task, as it effectively handles time-to-event data with extensive right-censoring—matching the characteristics of our dataset. In this section, we evaluate the utility of our proposed features.
Suitability of Survival Analysis.
Survival analysis is particularly apt for this problem because: (1) OSS projects may not all undergo cessation within the observation window, resulting in right-censored data. (2) Survival models naturally support event-time analysis, enabling either risk ranking (“which projects are likely to cessation sooner?”) or direct estimation of maintenance cessation time.
Rather than predicting a point estimate of cessation time, survival models rank project risk according to concordance with observed lifespans. The concordance index (C-index) is therefore the primary evaluation metric, measuring the fraction of predicted risk orderings that match actual event orderings. It ranges from 0.5 (random guessing) to 1.0 (perfect discrimination). Empirically, a C-index above 0.8 is considered strong for prediction (Liao et al., 2023).
Accelerated Failure Time (AFT) Model.
We instantiate the AFT approach using XGBoost’s survival module (Cox, 2018; Avinash Barnwal and Hocking, 2022), where the log of event time (maintenance lifespan) is regressed on input features.
The AFT model supports efficient optimization with right-censored samples using negative log-likelihood loss (nloglik). We partition the data into 80% training and 20% test splits, with early stopping and hyperparameter. We observe rapid convergence and stable loss minimization of the AFT (XGBoost-based) model in training.
Ablation Study and Predictive Contribution.
To rigorously assess the marginal and joint contribution of different feature groups, we conduct ablation experiments using AFT models trained on various combinations. Table 3 reports the resulting Harrel’s and Uno’s C-index scores for the following feature sets:
  • S: Surface features (in Section 4.1.1);
  • H: HITS influence scores (excluding design in Section 4.1.2);
  • U: User-centric features (in Section 4.1.2);
  • M: Maintainer-centric features (in Section 4.1.3);
  • P: Project evolution features (in Section 4.1.4).
Table 3. C-index of AFT models under different features.
Features Harrel’s C-index Uno’s C-index
S 0.748 0.653
S -- stars 0.689 0.621
S $+$ H 0.810 0.716
S $+$ U 0.826 0.751
S $+$ M 0.778 0.670
S $+$ P 0.767 0.665
U $+$ M $+$ P 0.838 0.773
All 0.846 0.781
We observe that augmenting the surface features (B) with any individual feature group consistently improves predictive performance. Notably, user-centric features (U) yield the largest gain, increasing Harrel’s C-index from 0.748 to 0.826—surpassing the effect of adding HITS scores (H) alone (0.810). Furthermore, using only U, M, and P—excluding all surface features—achieves near-parity with the full model (0.838 vs. 0.846), demonstrating the sufficiency and complementarity of the proposed features.
Feature Importance via F-score Analysis.
[…]
Figure 4. Feature importance across categories.
The results show that User-Centric Features contribute most prominently to the model, highlighting the predictive value of user engagement patterns (Ait et al., 2022). Maintainer-Centric Features follow with a substantial score of 2213.24, suggesting the importance of maintainer activity signals. Project Evolution Features also exhibit meaningful predictive strength, capturing long-term development dynamics. In contrast, Surface Features contribute the least, indicating limited standalone utility in cessation risk modeling.
Model-Agnostic Interpretability Using SHAP
To comprehensively validate feature influence, we employ SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017), a advanced method in explainable AI research (Borisov et al., 2021). SHAP quantifies the marginal contribution of each feature to predicted risk scores by averaging over all possible feature combinations, ensuring mathematically consistent and fair importance attribution. SHAP’s inherent robustness to feature collinearity ensures reliable importance attribution even when features exhibit moderate correlations.
The beeswarm SHAP visualization demonstrates that latest_ maintainer_activity_interval, weight_zscore, and contrib_diversity emerge as the three most influential predictors. These 3 features representing each of our novel feature categories, consistently exhibit the strongest directional impact on model predictions, confirming their critical role in cessation predicting.
Case Illustration: Temporal Dynamics of Feature Signals.
To illustrate the advantages of our newly engineered features, we examine the time-series evolution of multiple signals from the Electronic WeChat project (Zhongyi Tong et al., [n. d.]) (Figure 5). All feature values are min-max normalized to the [0.1, 1] range for a beautiful view.
Figure 5. Time-series feature dynamics for Electronic WeChat.
This project, a popular third-party MacOS WeChat client, ceased maintenance after Tencent officially released WeChat for Mac 2.0 on August 16, 2016 and updates regularly thereafter.
From Figure 5, we observe:
Surface Feature (Orange): The number of stars remained consistently high before the cessation event, reflecting the project’s lasting popularity. However, this persistent star growth results mainly from historical reputation and user recommendations, making it insensitive to changes in actual maintenance or activity. This indicates star count does not provide a timely warning of approaching maintenance cessation.
User-centric Feature (Blue): The percentile rank of interaction weight exhibits a steady rise as cessation approaches. This increase indicates a decline in genuine community engagement, even while stars remain high. As fewer influential users interact meaningfully with the project, this feature effectively signals a loss of active user support, often preceding maintenance termination.
Maintainer-centric Feature (Green): The time since the last maintainer activity increases continuously before maintenance cessation, showing that maintainers are gradually reducing their participation. This growing inactivity interval quantitatively foreshadows the decline of maintainer commitment, aligning with the actual cessation event.
Project Evolution Feature (Red): Initially, the maintainer contribution ratio is high but then drops sharply before cessation. This shift reflects a breakdown in the stable collaboration pattern, signaling impending inactivity.
This case underscores the critical need for advanced, multi-perspective features—beyond traditional surface features—when predicting OSS maintenance cessation with precision and interpretability. Our three newly proposed feature types capture a broader range of repository state information than surface activity features.
5. Real-World Deployment
In this section, we evaluate the practical utility of our maintenance cessation prediction system by deploying it within the openEuler ecosystem, a production-grade, community-driven Linux distribution, demonstrating both methodological generalizability and real-world impact.
5.1. Practical Risk Prediction via GBSA Classifier
To address the needs of stakeholders requiring binary, near-term risk assessments—for example, “Will this package likely cease maintenance within the next 6 months?”—we employ a Gradient Boosted Survival Analysis (GBSA) classifier using the feature set and data pipeline presented earlier. Unlike traditional regression-based survival models, GBSA excels at handling high-dimensional, time-varying, and right-censored datasets in a classification setting, and is well suited for integration into real-time package quality monitoring systems.
Formulation.
For deployment, we treat maintenance cessation prediction as a temporal binary classification problem. Let $T$ denote a reference date (in this study, July 1, 2018), and $\Delta t$ the prediction horizon (set to 6 months). For each repository, we aggregate all features up to $T$, and label as “positive” those repositories for which unambiguous maintenance cessation occurs between $T$ and $T+\Delta t$. Repositories not ceasing maintenance in this window are used as “negative” or censored samples. We select July 1, 2018 as the reference date $T$ because a relatively large number of openEuler ecosystem packages experienced maintenance cessation during this period of time. This maximizes the number and diversity of positive samples for robust model evaluation.
Training and Evaluation.
We balance positive and negative samples to address class imbalance and apply Bayesian optimization (via Optuna (Akiba et al., 2019)) for hyperparameter tuning (e.g., learning rate, tree depth, subsample ratio). Model performance is evaluated using accuracy, precision, recall, and the Harrell’s C-index, consistent with operational needs.
Our GBSA classifier achieves an overall prediction accuracy of 78.91% on an unbalanced held-out test set, confirming that our engineered features and survival-based modeling retain high discriminative power in a practical early warning context.
5.2. Deployment in the openEuler Ecosystem
openEuler (Zhou et al., 2022; openEuler, 2021) is a major open-source, Linux-based operating system, incubated within the OpenAtom Foundation and serving as a backbone for digital infrastructure in industries including Internet, finance, and telecommunications. In 2024 alone, openEuler registered over 5 million new installations (openEuler, 2024), reflecting significant exposure to the risks associated with unanticipated package maintenance cessation.
5.2.1. Mapping openEuler Packages to Upstream GitHub Repositories
openEuler is an open-source operating system whose source code is primarily hosted on the Gitee (Gitee, 2013), comprising two key repositories: the code repository for upstream source projects (openEuler, 2025a), and the package repository for build-ready software packages used in actual releases (openEuler, 2025c). These packages are directly tied to system distributions and user security. To assess ecosystem-level cessation risks, we take the openEuler release as a representative case and systematically map all its packages to their corresponding upstream GitHub repositories. Package-level metadata is retrieved from the official repository index (openEuler, 2025b).
Automated Mapping Pipeline.
We design a seven-stage pipeline to associate openEuler-22.03-LTS aarch64/loongarch64/x86_64 packages with their corresponding upstream GitHub repositories:
  • Rule-based mapping: Manually defined rules for common packages (e.g., texlive).
  • Gitee metadata: Extracting GitHub URLs from the url field in primary.xml.
  • Spec file parsing: Locating source repositories and extracting Source0 from .spec files via depchase (fedora-modularity et al., 2017).
  • YAML field matching: Parsing git_url or src_repo fields with typo tolerance.
  • URL-based inference: Inferring likely GitHub repos from homepage domains.
  • GitHub API search: Fallback search by name, ranking by stars ($\geq$32).
  • Homepage crawling: Scraping official websites to discover embedded GitHub links.
This automated pipeline successfully maps 14,284 out of 16,888 openEuler packages (coverage: 84.6%) to valid GitHub repositories, as visualized in Table 4. The resulting mapping forms the foundation for downstream risk inference.
Table 4. Distribution of mapping methods.
Method Proportion
Rule-based 59.32%
Gitee metadata 20.4%
Spec 13.7%
YAML 7.7%
URL-based 2.9%
API search 11.4%
Homepage crawling 2.4%
5.2.2. Maintenance Cessation Prediction: Experimental Protocol
For each mapped upstream repository, we extract all temporal features in Section 4.1, up to the reference time $T$, and apply the trained GBSA classifier to estimate the probability of maintenance cessation within the next 6 months.
Evaluation Setting.
[…]
Results.
Our deployed GBSA model successfully flags 3 out of 7 repositories that truly ceased maintenance within the package set of openEuler, based on a reference date of July 1, 2018 and a 6-month prediction window. This corresponds to multiple packages, with accuracy, precision, and C-index achieving 0.9857, 1.0, and 0.8049, respectively.
Interpretation.
A manual review of the four missed cases (false negatives) shows that these repositories lacked the typical warning signals captured by our features prior to cessation; most ceased maintenance abruptly or without clear early indicators. This highlights the limitation of behavioral features in detecting entirely unannounced or exogenous cessation events. Nevertheless, our model achieves strong overall accuracy and precision, reliably identifying early risk in the vast majority of practical cases, which supports deployment in real-world package management scenarios.
6. Threats to Validity
Internal Validity.
First, our definition of maintenance cessation, relying on explicit archival status or semantic cues in documentation, may still incompletely capture all true cessation events. Silent cessations, especially among small or low-profile projects, could be missed, while some ambiguous or template-based declarations might lead to mislabeling. Additionally, although our feature engineering seeks to comprehensively model user, maintainer, and project evolution dynamics, certain abrupt or exogenous cessation events (e.g., organizational, legal, or funding shocks) may not manifest detectable signals in repository behavior or metadata prior to cessation, potentially resulting in false negatives as observed in some openEuler cases (see Section 5.2). While we perform rigorous manual validation and sensitivity analyses, latent biases in label quality and feature coverage cannot be fully excluded.
External Validity.
Our study primarily targets GitHub-hosted repositories, which, while representing the majority of global OSS activity, may not generalize to other hosting platforms such as Gitee, GitLab, or self-hosted infrastructures. These alternatives often differ in developer conventions, community dynamics, and the richness or accessibility of behavioral data.
7. Conclusion
Maintenance cessation of open source software repositories poses major risks to software supply chains and ecosystem stability. In this work, we define a precise, actionable criterion for maintenance cessation, addressing the ambiguities of prior research. Building on this definition, we construct a comprehensive longitudinal dataset, supporting robust empirical analysis at scale. Our multi-perspective feature framework—integrating user influence, maintainer activity signals, and project evolution indicators—goes beyond traditional surface features, enabling more accurate prediction of maintenance cessation events. Using survival analysis, we achieve a high C-index (0.846), and confirmed the complementary value of our proposed features in different ways.
Finally, deployment within the openEuler ecosystem demonstrates the practical utility of our approach for proactive risk identification in real-world package management scenarios. Our interpretable modeling facilitates actionable interventions for OSS stakeholders, and paves the way for next-generation software supply chain risk management. As dependency on OSS continues to increase, scalable, early-warning tools such as ours are essential for ensuring the sustainability and security of the software ecosystem.